August 2010 Beam Test


In this logbook are reported some information needed for the Tracker shifters,

the complete loogbok of the testbeam is available at AMS E-Log Page.


Status of the data taking:

Point set Start Run Number End Run Number
Rich720 1281792481 1281824639
Tracker60FebBack 1281733073 1281755834
Tracker416_Rep01 1281703951 1281706488
Tracker10_Rep01 1281701726 1281702882
Tracker80_Rep01 1281694130 1281694130
Tracker60Feb 1281673346 1281693158
Tracker280 1281574137 1281668702
Tracker80 1281554781 1281571767
Tracker416 1281355854 1281550445
Tracker10 1281346173 1281354296
Tracker60 1281323720 1281340985
Ecal7 110 1281317427 1281321212


13/08/2010


  • 22:00 Starting TRACKER60FEBBACK good positions.
  • 16:00 AMS Rotation of 180 degrees around Y axis. The TTCS secondary FAC has been switched ON at a set point temperature of +24 degrees. The secondary pump has been switched on when the temperature difference between the liquid before the secondary pump and the accumulator was >3. After the rotation the primary system has been switched off. The temperature balance has been reached in few hours. The tracker temperatures and the TTCS parameters are OK.
  • 15:00 A bug in the decode has been fixed, now we have all the monitoring blocks on disk.


12/08/2010

  • 08:30 All the night calibrations checked. OK. Scanning restarted with much more prescaling, a dedicated calibration scanning started.


11/08/2010

  • Positions 250 to 262 - 4C622FBC -> 1281503164
  • Position 250 - 4C622E21 -> 1281502753
  • Positions 237 to 249 - 4C6222DC -> 1281499868
  • Position 237 - 4C62213B -> 1281499451
  • Positions 224 to 237 - 4C62165B -> 128149667
  • Position 224 - 4C61501 -> 1281496321
  • Positions 213 to 224 - 4C6208C9 -> 1281493193
  • Position 213 - 4C620755 -> 1281492821
  • Positions 203 to 212 - 4C61FBFE -> 1281489918
  • Position 203 - 4C61FA92 -> 1281489554
  • Positions 191 to 203 - 4C61EC37 -> 1281485879
  • Position 191 - 4C61EA5D -> 1281485405
  • Positions 178 to 191 - 4C61DD6B -> 1281482091
  • Position 178 - 4C61DC15
  • Positions 167 to 177 - 4C61CEB6 -> 1281478326
  • Position 167 - 4C61CD24 -> 1281477924
  • At 06:00, the delay is of 4 hours.
  • Decoding of BLOCKS has a two hour delay.
  • Run ID: 4C61BB95 (1281473429) for Position 154 Tracker 416 -> Whole run seems a failure. However Position 154 is present in Run 4C61BD2C.


10/08/2010

  • 11:40 - TTCS FAC heater was switched ON and GAC heater switched OFF under Johannes suggestion. However did not work and switched back to normal around 12:05.
  • 09:50 - T5-TDR2 Very High Noise on S-Side (CalDB_1281426490.root)
  • 09:49 - DAQ Started but still no beam. Cosmic Data taken.
  • 08:05 - AMS was powered OFF around 04:00 a.m. Latest News given by Mike is that the SPS is down because the kickers are broken. People from Cern are waiting 08:30 to access the SPS (waiting for low enough radiation) to evaluate the reparations. Until new info., V. Bindi and M. Heil switch ON AMS to take cosmics.


09/08/2010

  • 17:30 - more than 4000 strips masked.
  • Ladder High Occupancy Check
    * T0: TDR3 -> VA-13, TDR6,10,11,23 -> Few Strips, TDR19 -> VA-12-14.
    * T1: TDR3 -> Few Strips, TDR17 -> VA-1, TDR21 -> VA-12.
    * T2: TDR12 -> Good Clusters but with Very High Amplitude, TDR13 -> Few Strips
    * T3: TDR7 -> Different Gains on p and n and strange amplitude behaviour (ask to P. Azzarello), TDR14 -> VA-14-15.
    * T4: TDR10 -> Difference of Efficiency on p and n side because of some dead regions with no gain.
    * T6: TDR9,10,11 -> Few Strips, TDR13 -> VA-11, TDR22 -> VA-10-11-15.

CalDB _1281340982.root: problem with this calibration during Tracker60. In attachement the list of the bad ladders.


08/08/2010

  • Test Beam starting block: HRDL-B/0091/889
Topic attachments
I Attachment History Action Size Date Who Comment
PNGpng BeamProfileY_Prot180GeV.png r1 manage 9.5 K 2010-08-14 - 15:13 AlbertoOliva Y Beam Profile of secondary protons (180 GeV)
PNGpng BeamX.png r1 manage 10.0 K 2010-08-14 - 11:50 AlbertoOliva 400 GeV Protons Beam Profile
PNGpng BeamY.png r1 manage 10.2 K 2010-08-14 - 11:51 AlbertoOliva 400 GeV Protons Beam Profile
Texttxt CalDB_1281340982.txt r1 manage 0.2 K 2010-08-09 - 20:29 AlbertoOliva  
Texttxt TRACKER416_Ladders.txt r1 manage 23.3 K 2010-08-09 - 19:00 SadakazuHaino TkID/HwId of 416 positions
Texttxt TRACKER416_ams.txt r1 manage 17.1 K 2010-08-09 - 21:35 SadakazuHaino TRACKER416 positions in AMS coord. (x,y,z,dx,dy,dz)
PDFpdf TRACKER416_status.pdf r1 manage 759.6 K 2010-08-12 - 15:54 SadakazuHaino TRACKER416 statistics
Texttxt TRACKER60FEB_ams.txt r2 r1 manage 2.5 K 2010-08-13 - 10:39 SadakazuHaino TRACKER60 positions in AMS coord. (x,y,z,dx,dy,dz)
PDFpdf TRACKER60_status.pdf r1 manage 275.9 K 2010-08-14 - 06:47 SadakazuHaino TRACKER60 statistics
PDFpdf TRACKER80_z7_opt3.pdf r1 manage 13.3 K 2010-08-12 - 22:56 SadakazuHaino Missing 20 positions in TRACKER80
PNGpng TrackerEventSize_WithACC.png r1 manage 11.5 K 2010-08-14 - 15:12 AlbertoOliva Tracker Event Size Comparison (Beam FTorFTE - Beam FTandACC<1 - Cosmics)
PNGpng TrackerSize_Aug1020.png r1 manage 13.0 K 2010-08-14 - 11:53 AlbertoOliva Tracker Event Size Comparison (400 GeV Proton Beam - Cosmics)
PDFpdf rgt_4C6A41F1.pdf r1 manage 35.5 K 2010-08-17 - 16:27 SadakazuHaino Rigidity plot run 4C6A41F1 (60 GeV p/pi)
PDFpdf tb60feb_30.pdf r1 manage 20.9 K 2010-08-14 - 04:59 SadakazuHaino TRACKER60FEB position check 30
PDFpdf tbcomp_t60.pdf r1 manage 328.2 K 2010-08-13 - 19:13 SadakazuHaino Comparison of TRACKER60 positions between Aug. and Feb./2010
PDFpdf tbocp10.pdf r1 manage 28.5 K 2010-08-12 - 18:55 SadakazuHaino TRACKER10 track positions
PNGpng tbprof_60FEB.png r1 manage 8.0 K 2010-08-13 - 21:17 SadakazuHaino TRACKER60FEB positions and angles VS time
PNGpng tbprof_60FEBBACK.png r1 manage 7.5 K 2010-08-13 - 21:18 SadakazuHaino TRACKER60FEBBACK positions and angles VS time
PDFpdf tbtime_60feb.pdf r1 manage 15.9 K 2010-08-14 - 04:55 SadakazuHaino TRACKER60FEB statistics
PDFpdf tchk416_130_zr7.pdf r1 manage 22.5 K 2010-08-10 - 20:51 SadakazuHaino Beam profile check 416-130 (Zrot=-7cm) NEW TABLE
PDFpdf tchk416_27.pdf r1 manage 72.1 K 2010-08-09 - 18:06 SadakazuHaino Beam profile check 416-51 (Zrot=-15.2 cm)
PDFpdf tchk416_27_zr7.pdf r1 manage 20.0 K 2010-08-10 - 13:33 SadakazuHaino Beam profile check 416-27 (Zrot=-7cm)
PDFpdf tchk416_51_zr7.pdf r1 manage 19.3 K 2010-08-10 - 13:34 SadakazuHaino Beam profile check 416-51 (Zrot=-7 cm)
PDFpdf tchk416_71_zr7.pdf r1 manage 19.7 K 2010-08-10 - 13:43 SadakazuHaino Beam profile check 416-71 (Zrot=-7cm)
Edit | Attach | Watch | Print version | History: r32 < r31 < r30 < r29 < r28 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r32 - 2010-08-17 - SadakazuHaino
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback